### v1

@@ 1.2 Mitigating the impact of misleading visualizations
Adding a warning message provides non-significant changes of at most 4.9 percentage points, with most improvements observed on the real-world subset. This method assumes prior knowledge of the misleader present in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task [15, 16], as MLLMs struggle not only to distinguish non-misleading visualizations from misleading ones but also to correctly identify the specific misleader in a given visualization. Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of other correction methods, eliminating their negative impact on non-misleading visualizations.
We examine further the quality of the intermediate table extraction step using CHARTOM [12], which pairs each question with two visualizations of the same data — one misleading and one non-misleading. Ideally, the extracted tables should be identical for both. However, across all MLLMs, the extracted tables have a perfect match in only 4 out of 56 pairs (7.1%), and a partial match for 13 to 14 other pairs (23 to 25%). Notably, 3D effects achieve their deceptive purpose, as MLLMs systematically produce different tables when the visualization is presented with or without 3D effects. Since table-based QA and redrawing the visualization highly depend on this intermediate table extraction step, improving it is an important direction for future work.
@@ 2 Discussion
Our findings highlight the vulnerability of MLLMs to misleading visualizations, reducing their QA performance compared to non-misleading ones, down to the level of the random baseline. To address this vulnerability, we explored six inference-time correction methods. The most effective approach involves extracting the underlying table and performing the QA task using only the table, improving accuracy by up to 19.6 percentage points. While recent work in visualization comprehension with MLLMs primarily treats visualizations as images [17, 18, 19], our findings challenge this direction and underscore the merits of earlier approaches, such as DePlot [20], which use table extraction as an intermediate step, followed by table-based QA with a text-only model. By making MLLMs more robust to misleading visualizations, correction methods help ensure that AI-assisted chart reasoning does not yield inaccurate interpretations of the underlying data, reducing the risk that human users believe in and propagate misinformation.
We identify two limitations to our work. First, the visualization redrawing method, which uses the package Matplotlib, does not support maps, as Matplotlib lacks sufficient functionality for rendering high-quality maps. Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step.
@@ 3 Methods
3.1 Datasets
The misleading and non-misleading visualization datasets combine three existing resources and one introduced in this work. First, CALVI [8] includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ) with three to four choices. Second, CHARTOM [12] contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 ranking questions. The MCQs provide two to four choices. CHARTOM is the only dataset where each question is linked to two visualizations — one misleading and one non-misleading. Like CALVI, the underlying data is synthetic. Third, VLAT [13], the reference dataset to evaluate the human comprehension of visualizations, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. MCQs have two to four choices. Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. The real-world visualizations come from a collection annotated with misleader labels [1], which inspired the synthetic examples in CALVI. We manually create MCQs, using those from CALVI and CHARTOM as inspiration. The motivation for creating this additional resource is that CALVI and CHARTOM rely both on synthetic data. By incorporating questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
We also report the performance of MLLMs on the test set of ChartQA [9], which contains 2500 real-world visualizations paired with free-text questions, of which half are human-written and the others are AI-generated.
The datasets cover together 11 chart types: line charts, area charts, stacked area charts, bar charts, stacked bar charts, histograms, pie charts, scatter plots, bubble charts, maps, treemaps.
For free-text questions where the expected answer is a number, we adopt the relaxed accuracy metric of ChartQA [9]. Specifically, a prediction is considered correct if it falls within a ±5% interval of the ground truth. For MCQs, we present the choices in their default order. In Table 2, we report the mean accuracy of the MLLMs over three different choice orders. The choices are shuffled using three randomly generated numbers as seeds (654, 114, 25). Standard deviations for the accuracy range from 0.0 to 4.06 percentage points.
@@ 3.2 Inference with (M)LLMs
We evaluate 11 open-weight MLLMs from the Llava-Next [21], Qwen2VL [22], Ovis-1.6 [23], and InternVL2.5 [24] families, ranging from 2 to 38 billion parameters. Additionally, we include two commercial models with undisclosed number of parameters, GPT-4 and GPT-4o [10], as well as three open-weight MLLMs specifically trained for visualization reasoning: ChartInstruction [17], TinyChart [18], and ChartGemma [19] ranging from 3 to 13 billion parameters. We use the Hugging Face transformers [25] implementation for all open-weight (M)LLMs and access GPT models via the Azure OpenAI API. To ensure deterministic outputs, we fix the random seed at 42, set the temperature to 0, and use a top-p value of 1. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in a zero-shot manner. For TinyChart, we report results using the Direct approach [18].
@@ 3.3 Implementation of correction methods
All prompts are in the code, which is provided as supplementary material. The significance of correction methods is assessed using the McNemar test (p $\leq$ 0.05) [26]. The p-values are reported in Table 3.
@@
Table-based QA: the extracted table is provided as input with the question and an instruction to answer it based on the table.
Redrawn visualization: the text-only LLM receives the extracted table and the chart type as text input and generates Python code to create a visualization using Matplotlib. The code is then executed to produce a new visualization. If the code compiles successfully, the newly generated visualization replaces the original one in the QA prompt; otherwise, the original visualization remains in use.
@@ Acknowledgements
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We want to express our gratitude to Germàn Ortiz, Manisha Venkat, and Max Glockner for their feedback on a draft of this work.
Appendix A Definition of misleaders
Table 1: The 17 types of misleader included in the misleading visualizations dataset, with their number of occurrences ($n$) and definitions [1, 8].
@@ Appendix A Definition of misleaders
Appendix B Impact of choices order in MCQs
Table 2: Mean accuracy and standard deviation (%) when shuffling MCQ choices with three different random seeds.
@@ Appendix B Impact of choices order in MCQs
Appendix C P-values of significance tests
Table 3: P-values for the McNemar test assessing the significance of the correction methods. Significant results ($\leq$0.05) are marked in bold.
Method Model Misleading p-value Non-misleading p-value
qwen2vl 1.00 1.00
Misleader warning ovis1.6 0.17 0.25
internvl2.5 1.00 1.00
qwen2vl 0.58 0.63
Axes ovis1.6 0.38 0.08
internvl2.5 0.65 0.10
qwen2vl 0.21 1.00
Table ovis1.6 0.40 0.04
internvl2.5 0.80 0.66
qwen2vl 0.17 1.00
Table + Axes ovis1.6 1.00 0.09
internvl2.5 1.00 1.00
qwen2vl 1e-4 0.14
Table-based QA ovis1.6 3e-3 0.07
internvl2.5 5e-5 0.34
qwen2vl 0.01 0.45
Redrawn visualization ovis1.6 2e-4 0.45
internvl2.5 0.68 0.33

### v3

@@ 1.4.3 Other correction methods are not effective
None of the remaining correction methods yield significant improvements. Including the extracted table in the prompt even results in a significant decrease in accuracy on non-misleading visualizations for one MLLM. While using the extracted table alone proves to be the most effective method, combining it with the original misleading visualization does not produce similar gains. This further suggests that removing or replacing the misleading visualization, rather than merely supplementing it with additional prompt input, is a key factor in the success of a correction method. Although many misleaders rely on axis manipulations, providing the extracted axes alone has little to no effect on performance. Similarly, adding a warning message produces only non-significant changes, with accuracy gains of at most 4.9 pp, mostly limited to the real-world subset. This warning-based approach assumes prior knowledge of the specific misleader in the visualization, making the reported results an upper bound on its effectiveness. In practice, a classifier needs to detect first the presence of misleaders, which remains a challenging task [34, 35]. Given the already low results obtained with ground-truth misleader labels, this correction method appears unpromising overall. However, training a highly accurate misleader detection model could enable the selective application of correction methods. This would offer two key advantages: (1) avoiding the application of correction methods to non-misleading visualizations, thereby eliminating any risk of negative impact; and (2) allowing to select the most suitable correction method based on the type of misleader.
@@ 2 Discussion
Our findings highlight the vulnerability of MLLMs to misleading visualizations. Experiment 1 demonstrated that MLLMs are unable to accurately interpret the underlying data when presented with misleading visualizations, reducing their QA accuracy, on average, to the level of a random baseline. Experiment 2 showed that MLLMs are particularly susceptible to truncated axes, often failing to provide consistent answers when shown misleading and non-misleading versions of the same data. Identifying this vulnerability fills a critical gap in automated chart understanding research. While MLLMs report strong performance on standard benchmarks such as ChartQA, they remain unprotected against misleading visualizations. This raises serious concerns about their reliability in real-world settings, especially given the potential for such vulnerabilities to be exploited by malicious actors to spread disinformation [4] in polarized political contexts [17] or during crises like the COVID-19 pandemic [36]. Addressing this blind spot is an urgent priority for future research in automated chart understanding.
To mitigate this issue, we proposed six inference-time correction methods, two of which, table-based QA and redrawing the visualization, demonstrated significant improvements. However, table-based QA comes at the cost of reduced accuracy on non-misleading visualizations. While most recent work in chart understanding treats visualizations purely as images [26, 27, 28], our findings challenge this direction and sheds a new light on alternative approaches that involve table extraction as an intermediate step [37, 38, 39].
Our correction methods, in particular redrawing the visualization, relate to prior work on visualization linters [40, 41, 30]: rule-based methods to detect and correct misleaders. While these tools proved useful to make human readers more robust to misleading visualizations [30], they can only operate with specific chart design tools such as Matplotlib [33, 40, 41] or a very limited set of chart and misleader types [30]. In contrast, our correction methods do not impose such restrictions.
Our findings are further supported by prior work [42] and parallel studies [43, 44, 45], which also report MLLM vulnerabilities to misleaders [42, 43, 44] and to simple design decisions like the color palette [45], reinforcing the urgency and relevance of this new research direction. Our study differentiates itself in several key ways. Prior work [42] only evaluated one MLLM, GPT4, on six misleading visualizations, providing initial hints but lacking depth to fully establish the presence of a vulnerability. Unlike parallel works, we include real-world misleading visualizations and show that vulnerability is somewhat lower in these cases, especially for larger models which may rely on parametric knowledge. Beyond identifying the vulnerability, we provide the first evaluation of six correction methods. Crucially, we examine the trade-off between improving performance on misleading visualizations and preserving accuracy on non-misleading ones, an aspect not addressed in other works [44].
We identify three limitations to our work. First, the visualization redrawing method does not support maps, as Matplotlib [33] lacks sufficient functionality for rendering high-quality maps. Second, we assume prior knowledge of the chart type (bar, line, …) for generating the prompts for axes extraction and visualization redrawing. This is a reasonable assumption, as the chart type can either be provided by a human user or accurately predicted by a classifier as a preprocessing step. Third, existing datasets do not provide an equal representation of all misleaders, with truncated and inverted axes being the two most represented categories. This imbalance is prevalent in CALVI [13] and present to a lower extent in the real-world cases [7].
@@ 3 Methods
3.1 QA datasets
The misleading and non-misleading visualization QA datasets used in Experiments 1 and 3 combine three existing resources and one introduced in this work.
First, CALVI [13] includes 45 misleading and 15 non-misleading visualizations based on synthetic data, each paired with a multiple-choice question (MCQ) with three to four choices. The dataset was initially designed to evaluate human readers.
Second, CHARTOM [14, 15] contains 56 samples, including 28 MCQs, 20 free-text questions, and 8 rank questions. The MCQs provide two to four choices. Each question is linked to two visualizations, one misleading and one non-misleading. Like CALVI, the underlying data is synthetic, and the test was originally designed to evaluate humans. We changed the name of certain misleaders in CHARTOM to align them with the terminology used in CALVI [7, 13].
Third, VLAT [29], the reference dataset to evaluate the human comprehension of visualizations, provides 12 non-misleading visualizations, each paired with three to seven MCQs, for a total of 53 instances. The visualizations are based on real-world data. MCQs have two to four choices.
Fourth, we introduce a dataset of 42 real-world misleading visualizations, each annotated with a MCQ with three to four choices. They come from a collection annotated with the misleaders affecting them [7], which inspired the synthetic examples in CALVI. We manually create MCQs, using as reference those from CALVI and CHARTOM sharing the same misleader. The motivation for creating this additional resource is that CALVI and CHARTOM rely both on synthetic data. By incorporating questions about real-world data, we introduce direct conflicts with MLLMs’ parametric knowledge, allowing us to assess their vulnerability in real-world scenarios.
The non-misleading visualization dataset combines the synthetic instances of CALVI ($n$ = 15) of CHARTOM ($n$ = 56) with the real-world ones of VLAT ($n$ = 53).
We also provide the previously reported performance of MLLMs on the test set of ChartQA [18], which contains 2500 real-world visualizations paired with free-text questions, of which half are human-written and the others are AI-generated.
The datasets cover together 12 chart types: line charts, area charts, stacked area charts, bar charts, stacked bar charts, pictorial bars, histograms, pie charts, scatter plots, bubble charts, maps, treemaps.
For free-text questions where the expected answer is a number, we adopt the relaxed accuracy metric of ChartQA [18]. Specifically, a prediction is considered correct if it falls within a ±5% interval of the ground truth. For MCQs, we present the choices in their default order. In Table 4, we report the mean accuracy of the MLLMs over three different choice orders. The choices are shuffled using three randomly generated numbers as seeds (654, 114, 25). Standard deviations for the accuracy range from 0.0 to 4.06 pp.
3.2 Likert-scale dataset
For Experiment 2, we use four pairs of visualizations, one misleading and the other non-misleading, originally designed to evaluate humans [11]. The four pairs cover four unique chart types and three distinct misleaders.
@@ 3.3 Inference with (M)LLMs
All prompts are available in the code, which is provided as supplementary material. We use the Hugging Face transformers [46] implementation for all open-weight (M)LLMs and access GPT and Gemini models via the Azure OpenAI API and Google Cloud API, respectively. To ensure deterministic outputs, we set the random seed at 42, set the temperature to 0, and use a top-p value of 1. Following the standard ChartQA evaluation setup, all (M)LLMs are prompted in a zero-shot manner. For TinyChart, we report results using the Direct approach [27].
@@ 3.4 Implementation of correction methods
Table-based QA: the extracted table is provided as input in the prompt to the text-only LLM with the question and an instruction to answer it based on the table.
Redrawn visualization: the text-only LLM receives the extracted table and the chart type as text input and generates Python code to create a visualization using Matplotlib [33]. The code is then executed to produce a new visualization. If the code compiles successfully, the newly generated visualization replaces the original one in the QA prompt; otherwise, the original visualization remains in use.
[…]
Competing interests
The authors declare no competing interests.
[…]
Supplementary information
The datasets and code to reproduce the experiments are made available at github.com/UKPLab/arxiv2025-misleading-visualizations.
@@ Acknowledgements
This work has been funded by the LOEWE initiative (Hesse, Germany) within the emergenCITY center (Grant Number: LOEWE/1/12/519/03/05.001(0016)/72) and by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We want to express our gratitude to Niklas Traser for conducting an initial exploration of the real-world data, to Jan Zimny for our insightful discussions on the topic of misleading visualizations, and to Germàn Ortiz, Manisha Venkat, and Max Glockner for their feedback on a draft of this work.
Appendix A Extended data - Definition of misleaders
Table 1: Extended data - The 17 types of misleader included in this work, with their number of occurrences ($n$) and definitions [7, 13].
@@ Appendix A Extended data - Definition of misleaders
Appendix B Extended data - Datasets statistics
Table 2: Extended data - Datasets statistics. $n$ is the number of instances (question-answer-visualization triplets). $m$ is the number of misleader categories. MCQ indicates multiple-choie questions.
Dataset Experiments $n$ $m$ Question types
Misleading visualizations 1 & 3 143 17 MCQ, Free-text, rank
$\hookrightarrow$ CALVI (misleading) [13] 1 & 3 45 14 MCQ
$\hookrightarrow$ CHARTOM (misleading) [14, 15] 1 & 3 56 7 MCQ, Free-text, rank
$\hookrightarrow$ Real-world [7] 1 & 3 42 12 MCQ
Non-misleading visualizations 1 & 3 124 - MCQ, Free-text, rank
$\hookrightarrow$ CALVI (non-misleading) [13] 1 & 3 15 - MCQ
$\hookrightarrow$ CHARTOM (non-misleading) [14, 15] 1 & 3 56 - MCQ, Free-text, rank
$\hookrightarrow$ VLAT [29] 1 & 3 53 - MCQ
ChartQA [18] 1 2500 - Free-text
Lauer & O’Brien [11] 2 8 3 Likert-scale
Appendix C Extended data - Experiment 1 - Detailed accuracy scores
Figure 7: Extended data - Detailed accuracy scores (%) of Experiment 1 per misleader and MLLM.
Appendix D Extended data - Experiment 3 - significance tests
Table 3: Extended data - Test statistics (T) and P-values (p) for the McNemar test assessing the significance of the correction methods in Experiment 3. Significant results ($\leq$0.05) are marked in bold.
Method Model Misleading Non-misleading
T p T p
qwen2vl 5 1.00 0 1.00
Misleader warning ovis1.6 6 0.17 1 0.63
internvl2.5 4 1.00 0 1.00
qwen2vl 7 1.0 1 0.38
Axes ovis1.6 8 0.38 8 0.21
internvl2.5 5 1.00 5 1.00
qwen2vl 8 0.15 7 1.00
Table ovis1.6 9 0.40 11 0.04
internvl2.5 9 0.82 6 0.12
qwen2vl 6 0.12 6 0.45
Table + Axes ovis1.6 11 1.00 10 0.10
internvl2.5 8 1.00 10 0.68
qwen2vl 12 1e-4 10 0.05
Table-based QA ovis1.6 16 3e-3 10 0.05
internvl2.5 12 1e-4 17 0.43
qwen2vl 5 0.01 2 0.11
Redrawing visualization ovis1.6 5 2e-4 6 0.33
internvl2.5 11 0.84 6 0.24
Appendix E Extended data - Example of table extraction
Figure 8: Extended data - Table extraction applied to a non-misleading instance of CHARTOM and its misleading version with an inverted axis. Despite that the two visualizations represent the same data, the extracted tables show different values and trends.
Appendix F Extended data - Experiment 1 - Impact of choices order in MCQs
Table 4: Extended data - Mean accuracy (%) and standard deviation when shuffling MCQ choices of Experiment 1 with three different random seeds.